OpenStudy (anonymous):

find the limit of ln(tan x) as x approaches pi/2 from the negative side

OpenStudy (luigi0210):

Hello @maeyer0409 Welcome to Openstudy :)

OpenStudy (luigi0210):

Is this your question: \[\lim_{x \rightarrow \frac{\pi}{2}} \ln(tanx)\]

OpenStudy (anonymous):

yes but x as it approches from the negative side

OpenStudy (luigi0210):

Do you know how the graph looks like?

OpenStudy (anonymous):

no

OpenStudy (luigi0210):

It kind of looks like this: